### v2

# FIAN/TD/12-2024

Abstract
In this note we construct a formulation of $4d$ pure Yang–Mills theory in the unfolded language of higher-spin gravity. It represents a first-order formulation possessing manifest diffeomorphism- and gauge-invariance. We make use of an unfolding method that was recently put forward and used to unfold scalar electrodynamics. We also discuss features of various unfolding maps defined by the unfolded equations.
Introduction
Symmetry considerations are among the most important guiding principles in constructing theories of fundamental interactions. Higher-spin gravity, representing a theory of interacting massless fields of all spins, sets an example. To a large extent, this theory is determined by an infinite-dimensional higher-spin gauge symmetry [1]. For a partial review of the literature on the subject see [2].
To keep the higher-spin symmetry under control is of crucial importance. Originally, higher-spin gravity was formulated in the form of Vasiliev equations [3, 4] which are classical equations of motion written in the so-called unfolded form. Unfolded equations are first-order differential equations on unfolded fields being exterior forms, which possess manifest gauge symmetry. Thus, the unfolded dynamics approach [5, 3, 4, 6, 7, 8] represents a classical first-order formalism providing manifest diffeomorphism- and gauge-invariance. It also allows for an effective control over degrees of freedom, since unfolded fields (whose spectrum is infinite in dynamical field theories) parameterize all d.o.f. of a theory under consideration. In particular, this makes the unfolded dynamics approach a promising tool for the study of $AdS/CFT$-correspondence and other dualities [9, 10, 11]. A way to quantize classical unfolded theories was put forward in [12].
Apart from higher-spin gravity models [3, 4, 13, 14, 15, 16], not that many examples of unfolded formulations of nonlinear theories are available [17, 12, 18], which limits the ability to explore the full potential of the approach. The reason behind this was a lack of a practical algorithm of unfolding, which was especially critical for nonlinear theories. Therefore, linear models were mainly studied [19, 20, 21, 22, 23]. In [18], a novel unfolding method was put forward, which allowed one to construct an unfolded formulation of $4d$ scalar electrodynamics. In brief, it consists in the initial postulation of the form of an unfolded master-field and the subsequent derivation of unfolded equations as identities for this master-field.
In this note, we apply the method of [18], slightly improving it, in order to construct an unfolded formulation of $4d$ on-shell pure Yang–Mills theory.
The note is organized as follows. First, we present basic definitions of the unfolded dynamics approach. Next, we construct unfolded Yang–Mills equations by fixing the form of the unfolded master-fields and using certain operator relations that we derive. Then we discuss the properties of the resulting unfolded system and comment on the unfolding maps it defines. In conclusion, we provide a brief overview of possible directions for further research.
Unfolded dynamics approach
Within the unfolded dynamics approach [5, 3, 4, 6, 7, 8], a field theory is formulated in terms of first-order equations
$\mathrm{d}W^{A}(x)+G^{A}(W)=0$ (1)
on unfolded fields $W^{A}(x)$, which are exterior forms on a space-time manifold $M^{d}$. Here $A$ stands for all indices of an unfolded field, $\mathrm{d}$ is the exterior differential on $M^{d}$ and $G^{A}(W)$ is built from exterior products of unfolded fields (the wedge symbol is omitted throughout the paper). For every $W^{A}$ there is one and only one unfolded equation (1).
There is a consistency condition for the system (1)
$G^{B}\dfrac{\delta G^{A}}{\delta W^{B}}\equiv 0,$ (2)
which follows from $\mathrm{d}^{2}\equiv 0$ and restricts possible $G^{A}$. Considering this, unfolded equations (1) possess a manifest gauge symmetry [^1] Strictly speaking, this is true provided that consistency holds regardless of the dimension of $M^{d}$, which is the case for all known examples [8].
$\delta W^{A}=\mathrm{d}\varepsilon^{A}(x)-\varepsilon^{B}\dfrac{\delta G^{A}}{\delta W^{B}},$ (3)
so that every $(n>0)$-form field $W^{A}$ gives rise to a gauge symmetry with a $(n-1)$-form parameter $\varepsilon^{A}(x)$. 0-form fields are transformed only under gauge symmetries of 1-form fields via the second term in (3).
In dynamical field theories, a spectrum of unfolded fields is infinite, since they parameterize all d.o.f. Usually, the space of unfolded fields possesses some grading bounded from below, so that equations (1) express (perhaps, in a very complicated nonlinear way) higher-grade fields in terms of derivatives of the lower-grade ones. In addition, (1) may implicitly impose dynamical constraints on the lowest-grade fields, in which case the system is said to be on-shell.
Ultimately, an unfolded system (1) describes a theory of lowest-grade fields (also referred to as primaries), possibly subjected to some differential constraints (e.o.m. or whatever), while higher-grade fields form (infinite) towers of their covariant differential descendants. Importantly, an unfolded formulation is manifestly diffeomorphism- and gauge-invariant and provides the control over d.o.f. of the theory. All this makes unfolded dynamics approach a powerful tool for studying higher-spin gravity. But besides higher spins, these remarkable features can be useful for more standard field theories as well. Here our goal is to reformulate $4d$ pure Yang–Mills theory in the unfolded form (1), making use of the unfolding method proposed in [18].
Yang–Mills fields and auxiliary spinors
In $sl(2,\mathbb{C})$-spinor notations, $4d$ Yang–Mills equations together with Bianchi identities are
$\mathrm{D}_{\beta\dot{\alpha}}F^{\beta}{}_{\alpha}=0,\hskip 10.00002pt\mathrm{D}_{\alpha\dot{\beta}}\bar{F}^{\dot{\beta}}{}_{\dot{\alpha}}=0,$ (4)
where the (anti-)self-dual components of the strength tensor are
[…]
and the covariant derivative is
$\mathrm{D}_{\alpha\dot{\alpha}}:=\frac{\partial}{\partial x^{\alpha\dot{\alpha}}}-i[A_{\alpha\dot{\alpha}},\bullet],$ (6)
so that
$[\mathrm{D}_{\alpha\dot{\alpha}},\mathrm{D}_{\beta\dot{\beta}}]=-i\epsilon_{\alpha\beta}\bar{F}_{\dot{\alpha}\dot{\beta}}-i\epsilon_{\dot{\alpha}\dot{\beta}}F_{\alpha\beta}.$ (7)
Antisymmetric spinor metric
$\epsilon_{\alpha\beta}=\epsilon_{\dot{\alpha}\dot{\beta}}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),\quad\epsilon^{\alpha\beta}=\epsilon^{\dot{\alpha}\dot{\beta}}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$ (8)
moves spinor indices up/down according to
$v_{\alpha}=\epsilon_{\beta\alpha}v^{\beta},\quad v^{\alpha}=\epsilon^{\alpha\beta}v_{\beta},\quad\bar{v}_{\dot{\alpha}}=\epsilon_{\dot{\beta}\dot{\alpha}}\bar{v}^{\dot{\beta}},\quad\bar{v}^{\dot{\alpha}}=\epsilon^{\dot{\alpha}\dot{\beta}}\bar{v}_{\dot{\beta}}.$ (9)
Indices of multispinors, denoted with the same letter, are either contracted or symmetrized, depending on their relative positions (the same holds for dotted indices)
$T_{\alpha\alpha}:=T_{(\alpha_{1}\alpha_{2})},\quad T_{\alpha}{}^{\alpha}:=\epsilon^{\alpha\beta}T_{\alpha\beta}.$ (10)
In order to unfold Yang–Mills theory, one has to introduce, on top of primaries $F_{\alpha\alpha}$ and $\bar{F}_{\dot{\alpha}\dot{\alpha}}$, the infinite towers of all their differential on-shell descendants. This can be conveniently performed by means of auxiliary commuting spinors [^2] In $4d$ Vasiliev equations [3, 4], these spinors become generators of an associative higher-spin gauge algebra through a certain non-commutative star product defined on them. $Y=(y^{\alpha},\bar{y}^{\dot{\alpha}})$, then the whole towers get packed into unfolded Yang–Mills master-fields, which we postulate to be of the form
$F(Y|x)=e^{\mathrm{D}y\bar{y}}F_{\alpha\alpha}(x)y^{\alpha}y^{\alpha}e^{-\mathrm{D}y\bar{y}},\quad\bar{F}(Y|x)=e^{\mathrm{D}y\bar{y}}\bar{F}_{\dot{\alpha}\dot{\alpha}}(x)\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}e^{-\mathrm{D}y\bar{y}},$ (11)
where right-acting derivatives are contracted with spinors as
$\mathrm{D}y\bar{y}:=y^{\alpha}\bar{y}^{\dot{\alpha}}\mathrm{D}_{\alpha\dot{\alpha}}.$ (12)
Unfolded master-fields (11) contain primary Yang-Mills tensors as (anti-)holomorphic in $Y$ components
$F_{\alpha\alpha}(x)y^{\alpha}y^{\alpha}=F(Y|x)|_{\bar{y}=0},\quad\bar{F}_{\dot{\alpha}\dot{\alpha}}(x)\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}=\bar{F}(Y|x)|_{y=0},$ (13)
together with an infinite sequence of their fully symmetrized traceless covariant derivatives of all orders. This constitutes a set of all independent covariant descendants of the primary Yang–Mills tensor, since antisymmetrizations and contractions are determined by (4) and (7). By construction, unfolded master-fields (11) inherit the adjoint representation of the gauge algebra of primary tensors.
The aforementioned grading on the space of unfolded fields can be introduced in terms of the spinor Euler operators
$N:=y^{\alpha}\partial_{\alpha},\quad\bar{N}:=\bar{y}^{\dot{\alpha}}\bar{\partial}_{\dot{\alpha}},$ (14)
where $\partial_{\alpha}$ and ¯$\bar{\partial}_{\dot{\alpha}}$ are $y^{\alpha}$- and $\bar{y}^{\dot{\alpha}}$-derivatives. From their definitions (11), master-fields obey
$(N-\bar{N})F=2F,\quad(N-\bar{N})\bar{F}=-2\bar{F},$ (15)
which corresponds to helicities ±$\pm 1$ (in higher-spin gravity, strength tensors with $|N-\bar{N}|F=2sF$ describe spin-$s$ massless fields) and, further,
$\mathrm{D}y\bar{y}F=\bar{N}F,\quad\mathrm{D}y\bar{y}\bar{F}=N\bar{F}.$ (16)
Then a grading operator on the space of unfolded fields can be defined as
$\mathcal{G}:=\frac{1}{2}(N+\bar{N}-2).$ (17)
It decomposes the space of master-fields $F$ into a direct sum of its eigenspaces (the same applies to $\bar{F}$)
$F(Y|x)=\sum_{n=0}^{\infty}F^{(n)}(Y|x),\quad\mathcal{G}F^{(n)}=nF^{(n)},$ (18)
so that the primary has 00-grade, while $n$-grade unfolded fields represent its covariant derivatives of $n$-th order, as follows from (11).
To formulate the unfolded equations (1), one needs to express the derivatives of the unfolded fields in algebraic terms. In our particular case, the task is to express $\mathrm{D}_{\alpha\dot{\beta}}F$ for $F$ (11) in terms of $Y$ and $\partial/\partial Y$ acting on $F$ and $\bar{F}$ (and the same for $\mathrm{D}_{\alpha\dot{\beta}}\bar{F}$). Before we start processing (11), let us work out some general operator relations, which simplify the analysis implied by the method of [18]. Combining a commutator formula
[…]
with an Euler-operator representation of a homotopy integral
[…]
one obtains
$[\hat{\mathrm{A}},e^{\hat{\mathrm{D}}}]=(\frac{1}{N_{\hat{\mathrm{D}}}}e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}})e^{\hat{\mathrm{D}}}=e^{\hat{\mathrm{D}}}(\frac{1}{N_{\hat{\mathrm{D}}}}e^{-\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{\hat{\mathrm{D}}}),$ (21)
where the inverse Euler operator is understood as (for $F(0)=0$)
[…]
Applying this to an operator of the form
$\hat{\mathrm{B}}=e^{\hat{\mathrm{D}}}\hat{\mathrm{C}}e^{-\hat{\mathrm{D}}}$ (23)
yields
$[\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}}e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}.$ (24)
Next, considering the case $\hat{\mathrm{C}}=\hat{\mathrm{D}}$, from (24) one gets
$e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}=[\hat{\mathrm{A}},\hat{\mathrm{D}}]-\frac{1}{N_{\hat{\mathrm{D}}}-1}(e^{\hat{\mathrm{D}}}[[\hat{\mathrm{A}},\hat{\mathrm{D}}],\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}),$ (25)
so that (24) can be equivalently rewritten as
$[\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}(N_{\hat{\mathrm{D}}}-1)}e^{\hat{\mathrm{D}}}[\hat{\mathrm{D}},[\hat{\mathrm{A}},\hat{\mathrm{D}}]]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+[[\hat{\mathrm{A}},\hat{\mathrm{D}}],\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}.$ (26)
For convenience, we denote an "unfolding map" for an arbitrary function $\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)$ as
$\ll\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)\gg:=e^{\mathrm{D}y\bar{y}}\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)e^{-\mathrm{D}y\bar{y}},$ (27)
so that, in particular,
$F(Y|x)=\ll F_{\alpha\alpha}(x)y^{\alpha}y^{\alpha}\gg,\quad\bar{F}(Y|x)=\ll\bar{F}_{\dot{\alpha}\dot{\alpha}}(x)\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg.$ (28)
Then for an arbitrary $\mathrm{C}(Y|x)$ one has from (26)
$\ll\partial_{\mu}\mathrm{C}\gg=(\partial_{\mu}-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}})\ll\mathrm{C}\gg+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\ll\mathrm{C}\gg],$ (29)
¯$\ll\bar{\partial}_{\dot{\mu}}\mathrm{C}\gg=(\bar{\partial}_{\dot{\mu}}-\mathrm{D}_{\alpha\dot{\mu}}y^{\alpha})\ll\mathrm{C}\gg+i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+1)(\bar{N}+2)}F,\ll\mathrm{C}\gg].$ (30)
From here on, all derivatives inside the angle brackets always act only on an expression within brackets and never differentiate unfolding exponents of (27). The square brackets stand for the commutator in the gauge Lie algebra and the Euler operators of (26) are expressed in terms of spinor Euler operators (14). Master-fields $F$ and $\bar{F}$ arise in (29)-(30) through (7) without assuming Yang–Mills equations (4).
Now let us get down to solving the problem. Direct application of (24) to $\mathrm{D}_{\mu\dot{\mu}}F(Y|x)$ gives
$\mathrm{D}_{\mu\dot{\mu}}F=[\frac{1}{N}\ll[\mathrm{D}_{\mu\dot{\mu}},\mathrm{D}y\bar{y}]\gg,F]+\ll\mathrm{D}_{\mu\dot{\mu}}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg.$ (31)
The task is to get rid of all covariant derivatives and angle brackets on the r.h.s. by re-expressing them in terms of unfolded master-fields $F$ and $\bar{F}$ acted on by $Y$’s and $\frac{\partial}{\partial Y}$’s. To accomplish this, we have the following tools at our disposal: Yang–Mills equations (4) together with (7), relations (29) and (30), Schouten identities for spinors and Jacobi identity of the gauge Lie algebra.
Applying (7) to the first term on the r.h.s. of (31) and (4) (plus Schouten identities) to the second one, one has
$\mathrm{D}_{\mu\dot{\mu}}F=\frac{i}{2}\bar{y}_{\dot{\mu}}[\frac{1}{N}\ll\partial_{\mu}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg,F]+\frac{i}{2}y_{\mu}[\frac{1}{\bar{N}}\ll\bar{\partial}_{\dot{\mu}}\bar{F}_{\dot{\alpha}\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg,F]+\frac{1}{3}\ll\partial_{\mu}\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg.$ (32)
First, we process the first term on the r.h.s. (the second term will be resolved by conjugation). Applying (29) to the problematic factor yields
$\ll\partial_{\mu}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\partial_{\mu}F-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F].$ (33)
So, one needs to process $\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F$. Contracting (32) with $\bar{y}^{\dot{\mu}}$ gives
$\mathrm{D}_{\mu\dot{\mu}}\bar{y}^{\dot{\mu}}F=iy_{\mu}[\frac{1}{\bar{N}-1}\bar{F},F]+\frac{1}{3}\ll\partial_{\mu}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg.$ (34)
On the other hand, from (16) and (29) one finds
$\ll\partial_{\mu}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\partial_{\mu}\bar{N}F-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{N}F+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\bar{N}F].$ (35)
Combining (34) and (35), one obtains after rearranging the Euler-operator ratios
$\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F=\frac{\bar{N}}{N+1}\partial_{\mu}F+iy_{\mu}\frac{2}{N+2}[\frac{1}{N+2}\bar{F},F]+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F].$ (36)
Thus one finds for (33)
$\ll\partial_{\mu}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\frac{2}{N+1}(\partial_{\mu}F-iy_{\mu}[\frac{1}{N+2}\bar{F},F]).$ (37)
Conjugation gives
¯$\ll\bar{\partial}_{\dot{\mu}}\bar{F}_{\dot{\alpha}\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg=\frac{2}{\bar{N}+1}(\bar{\partial}_{\dot{\mu}}\bar{F}-i\bar{y}_{\dot{\mu}}[\frac{1}{\bar{N}+1}F,\bar{F}]).$ (38)
These bring first two terms on the r.h.s. of (32) to the admissible form.
Now we process the last term in (32). By virtue of (29) one has
$\ll\partial_{\mu}\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=(\partial_{\mu}-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}})\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg+$
$+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg].$ (39)
Since obviously
¯$\ll\bar{\partial}_{\dot{\mu}}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=0,$ (40)
one has from (30)
$\mathrm{D}_{\alpha\dot{\mu}}y^{\alpha}F=\bar{\partial}_{\dot{\mu}}F+i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+1)(\bar{N}+2)}F,F].$ (41)
Using this and contracting (32) with $y^{\mu}$ yields
¯$\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\bar{\partial}_{\dot{\mu}}F-i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+2)}F,F],$ (42)
which together with (36) turns (39) to
$\ll\partial_{\mu}\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\partial_{\mu}\bar{\partial}_{\dot{\mu}}F-i\bar{y}_{\dot{\mu}}\partial_{\mu}[\frac{1}{\bar{N}+2}F,F]+i[\frac{1}{(N+1)(N+2)}\bar{F},y_{\mu}\bar{\partial}_{\dot{\mu}}F]+$
$+y_{\mu}\bar{y}_{\dot{\mu}}[\frac{1}{(N+1)(N+2)}\bar{F},[\frac{1}{\bar{N}+2}F,F]]+\mathrm{D}_{\mu\dot{\mu}}F-\frac{\bar{N}+1}{N+1}\partial_{\mu}\bar{\partial}_{\dot{\mu}}F-$
$-2iy_{\mu}\bar{\partial}_{\dot{\mu}}\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F]-iy_{\mu}\bar{\partial}_{\dot{\mu}}[\frac{1}{(N+1)(N+2)}\bar{F},F]+$
$+i\bar{y}_{\dot{\mu}}[\frac{1}{\bar{N}+1}(\frac{\bar{N}}{N+1}\partial_{\mu}F+iy_{\mu}\frac{2}{N+2}[\frac{1}{N+2}\bar{F},F]+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F]),F]+$
$+i\bar{y}_{\dot{\mu}}[\frac{1}{N}F,(\frac{\bar{N}}{N+1}\partial_{\mu}F+iy_{\mu}\frac{2}{N+2}[\frac{1}{N+2}\bar{F},F]+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F])].$ (43)
Substituting (37), (38) and (43) into (32) and combining like terms using the gauge algebra Jacobi identity, the final result is
$\mathrm{D}_{\mu\dot{\mu}}F=\frac{1}{N+1}\partial_{\mu}\bar{\partial}_{\dot{\mu}}F+iN[\frac{1}{N(N+1)}\bar{y}_{\dot{\mu}}\partial_{\mu}F,\frac{1}{N}F]-iy_{\mu}\bar{\partial}_{\dot{\mu}}\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F]+$
$+[\frac{i}{(N+1)(N+2)}y_{\mu}\bar{\partial}_{\dot{\mu}}\bar{F},F]+\frac{1}{2}y_{\mu}\bar{y}_{\dot{\mu}}[\frac{N+3}{(N+1)(N+2)}[\frac{1}{N+2}\bar{F},F],F]+$
$+\frac{3}{2}y_{\mu}\bar{y}_{\dot{\mu}}[\frac{1}{(N+1)(N+2)}[\frac{1}{N}F,\bar{F}],F]+y_{\mu}\bar{y}_{\dot{\mu}}[\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F],\frac{1}{N}F].$ (44)
Poincaré symmetry and diffeomorphism-invariance
The manifestly Poincaré-invariant relation (44) is written in Cartesian coordinates, together with (6) and (11), since they involve $\frac{\partial}{\partial x^{\mu\dot{\mu}}}$ inside of $\mathrm{D}_{\mu\dot{\mu}}$. The unfolded dynamics approach requires manifest coordinate-independence, which is ensured by formulating equations in terms of exterior forms. To attain this, we switch to the fiber space picture: we claim that $F(Y|x)$ and $\bar{F}(Y|x)$ are now 0-forms on the Minkowski base manifold with some local coordinates $x^{\underline{n}}$, while spinor variables $\{y^{\alpha},\bar{y}^{\dot{\alpha}}\}$ are coordinates in the fiber. This requires appropriate generalization of $\mathrm{D}_{\mu\dot{\mu}}$ in (44).
According to the ideology of unfolding, global Poincaré symmetry of a relativistic theory should arise in terms of the unique general formula (3). This is achieved by introducing a non-dynamical 1-form $\Omega(x)$, which takes values in Lie algebra $iso(1,3)$
$\Omega=e^{\alpha\dot{\beta}}P_{\alpha\dot{\beta}}+\omega^{\alpha\alpha}M_{\alpha\alpha}+\bar{\omega}^{\dot{\alpha}\dot{\alpha}}\bar{M}_{\dot{\alpha}\dot{\alpha}},$ (45)
with $P_{\alpha\dot{\alpha}}$, $M_{\alpha\alpha}$ and $\bar{M}_{\dot{\alpha}\dot{\alpha}}$ being generators of translations and rotations of $\mathbb{R}^{1,3}$, and $e^{\alpha\dot{\beta}}$ and $\omega^{\alpha\alpha}$ ($\bar{\omega}^{\dot{\alpha}\dot{\alpha}}$) being 1-forms of a vierbein and a Lorentz connection.
$\Omega$ is subjected to the flatness condition (square brackets stand for the $iso(1,3)$-commutator)
$\mathrm{d}\Omega+\frac{1}{2}[\Omega,\Omega]=0,$ (46)
so that the corresponding gauge symmetry (3) is
$\delta\Omega=\mathrm{d}\varepsilon(x)+[\Omega,\varepsilon]$ (47)
and describes an infinite-dimensional freedom in switching between all possible local coordinates on $\mathbb{R}^{1,3}$. This boils down to 10-dimensional global Poincaré symmetry after fixing some particular solution $\Omega_{0}$ and restricting to those residual $\varepsilon(x)$ which leave it invariant
$\mathrm{d}\varepsilon_{0}+[\Omega_{0},\varepsilon_{0}]=0.$ (48)
The simplest non-degenerate global solution to (46) is provided by Cartesian coordinates
$e_{\underline{m}}{}^{\alpha\dot{\beta}}=(\bar{\sigma}_{\underline{m}})^{\dot{\beta}\alpha},\quad\omega_{\underline{m}}{}^{\alpha\alpha}=0,\quad\bar{\omega}_{\underline{m}}{}^{\dot{\alpha}\dot{\alpha}}=0,$ (49)
with global symmetries (48) parameterized by $x$-independent $\xi^{\alpha\dot{\beta}}$, $\xi^{\alpha\alpha}$ and $\bar{\xi}^{\dot{\alpha}\dot{\alpha}}$
$\varepsilon_{0}^{\alpha\dot{\beta}}=\xi^{\alpha\dot{\beta}}+\xi^{\alpha}\text{}_{\gamma}(\bar{\sigma}_{\underline{m}})^{\dot{\beta}\gamma}x^{\underline{m}}+\bar{\xi}^{\dot{\beta}}\text{}_{\dot{\gamma}}(\bar{\sigma}_{\underline{m}})^{\dot{\gamma}\alpha}x^{\underline{m}},\quad\varepsilon_{0}^{\alpha\alpha}=\xi^{\alpha\alpha},\quad\bar{\varepsilon}_{0}^{\dot{\alpha}\dot{\alpha}}=\bar{\xi}^{\dot{\alpha}\dot{\alpha}}.$ (50)
Analogously, in order to realize the Yang–Mills gauge symmetry via (3), one introduces a 1-form $A(x)$, with $A_{\alpha\dot{\alpha}}(x)$ being its expansion in the vierbein
$A(x)=e^{\alpha\dot{\alpha}}A_{\alpha\dot{\alpha}}.$ (51)
Now an appropriate coordinate-independent generalization of $\mathrm{D}_{\mu\dot{\mu}}$ is a 1-form operator $\mathrm{D}$, supplemented by Lorentz-connection terms rotating fiber coordinates $Y$,
$\mathrm{D}:=\mathrm{d}+\omega^{\alpha\alpha}y_{\alpha}\partial_{\alpha}+\bar{\omega}^{\dot{\alpha}\dot{\alpha}}\bar{y}_{\dot{\alpha}}\bar{\partial}_{\dot{\alpha}}-i[A,\bullet].$ (52)
In Cartesian coordinates (49), this indeed boils down to
$\mathrm{D}=\mathrm{d}x^{\mu\dot{\mu}}\mathrm{D}_{\mu\dot{\mu}}.$ (53)
Unfolded Yang–Mills equations and unfolding maps
Now we are ready to write down an unfolded system for Yang–Mills theory. Contracting (5) and (44) with the vierbeins yields
$\mathrm{d}A+[A,A]=\frac{1}{4}e^{\alpha}{}_{\dot{\beta}}e^{\alpha\dot{\beta}}\partial_{\alpha}\partial_{\alpha}F|_{\bar{y}=0}+\frac{1}{4}e_{\beta}{}^{\dot{\alpha}}e^{\beta\dot{\alpha}}\bar{\partial}_{\dot{\alpha}}\bar{\partial}_{\dot{\alpha}}\bar{F}|_{y=0},$ (54)
$\mathrm{D}F=\frac{1}{N+1}e\partial\bar{\partial}F+iN[\frac{1}{N(N+1)}e\partial\bar{y}F,\frac{1}{N}F]-iey\bar{\partial}\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F]+$
$+[\frac{i}{(N+1)(N+2)}ey\bar{\partial}\bar{F},F]+\frac{1}{2}ey\bar{y}[\frac{N+3}{(N+1)(N+2)}[\frac{1}{N+2}\bar{F},F],F]+$
$+\frac{3}{2}ey\bar{y}[\frac{1}{(N+1)(N+2)}[\frac{1}{N}F,\bar{F}],F]+ey\bar{y}[\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F],\frac{1}{N}F]$ (55)
plus a conjugate equation for $\bar{F}$ resulting from exchanging barred and unbarred objects in (55). Here
$e\partial\bar{\partial}:=e^{\alpha\dot{\beta}}\partial_{\alpha}\bar{\partial}_{\dot{\beta}},\quad e\partial\bar{y}:=e^{\alpha\dot{\beta}}\partial_{\alpha}\bar{y}_{\dot{\beta}},\quad ey\bar{\partial}:=e^{\alpha\dot{\beta}}y_{\alpha}\bar{\partial}_{\dot{\beta}}\quad ey\bar{y}:=e^{\alpha\dot{\beta}}y_{\alpha}\bar{y}_{\dot{\beta}}.$ (56)
A full spectrum of unfolded fields consists of a 1-form $\Omega$ describing Minkowski background, a 1-form of the gauge potential $A$ and the 0-forms of master-fields $F(Y|x)$ and $\bar{F}(Y|x)$ encoding the Yang–Mills tensor together with an infinite tower of its covariant derivatives. Corresponding unfolded equations are (46), (54) and (55) plus a conjugate for $\bar{F}$. The formulation is manifestly diffeomorphism-invariant. 1-forms $\Omega$ and $A$ give rise to two manifest symmetries in accordance with (3): the global (after fixing $\Omega$) Poincaré one and the local Yang–Mills one. The Yang–Mills symmetry is realized as
$\delta A(x)=\mathrm{D}\varepsilon(x),\quad\delta F(Y|x)=i[\varepsilon(x),F(Y|x)],\quad\delta\bar{F}(Y|x)=i[\varepsilon(x),\bar{F}(Y|x)].$ (57)
By construction, the unfolded system is consistent, provided unfolded 0-form fields obey the helicity constraint (15). At the same time, a direct check of (2) seems hardly executable due to the complexity of the equations. One of the possible forms of solution to the system is (11) with primary fields subjected to (4). Let us quickly derive this directly from (55).
Master-fields $F$ and $\bar{F}$ are assumed to be analytical in $Y$ and obey (15). For the sake of brevity, here and below we consider only an anti-selfdual component $F$, but everything applies to $\bar{F}$ as well. Acting on (55) with $y^{\beta}\bar{y}^{\dot{\beta}}\frac{\delta}{\delta e^{\beta\dot{\beta}}}$ produces (16), whose solution, accounting for (15), is (11). On the other hand, acting on (55) with $\frac{1}{2}\partial_{\mu}\partial^{\beta}\frac{\delta}{\delta e^{\beta\dot{\mu}}}$ and putting $\bar{y}^{\dot{\alpha}}=0$ yields, accounting for (15) again,
$\mathrm{D}_{\beta\dot{\mu}}F^{\beta}{}_{\mu}=0.$ (58)
Thus, the unfolded system presented above indeed provides a consistent manifestly diffeomorphism- and gauge-invariant first-order formulation of $4d$ Yang-Mills theory. Unfolded master-fields $F(Y|x)$ and $\bar{F}(Y|x)$ encode all on-shell d.o.f. of the Yang-Mills tensor as expansions in auxiliary spinors $Y$.
The system allows for two obvious reductions. The first one is the anti-selfdual case $\bar{F}(Y|x)=0$ (or the selfdual $F(Y|X)=0$), then only first two terms on the r.h.s. of (55) survive. In different (tensor) terms, this was presented in [24]. The second one is the abelian case with all commutators vanishing, then only the first term on the r.h.s. of (55) remains.
One can think of the equation (55) as defining an unfolding map from $x$-space to $Y$-space
$F_{\alpha\alpha}(x)|_{on-shell}\rightarrow F(Y|x)\rightarrow\mathcal{F}(Y):=F(Y|x=0).$ (59)
Then (11) explicitly realizes the first arrow in (59). The field $\mathcal{F}(Y)$ carries precisely the same information as on-shell $F_{\alpha\alpha}$ does. In a sense, spinors $Y=\{y^{\alpha};\bar{y}^{\dot{\alpha}}\}$ effectively replace space-time coordinates $x^{\alpha\dot{\alpha}}$ for on-shell configurations, hence being conjugate to spinor-helicity variables that resolve light-like momenta $p^{\alpha\dot{\alpha}}=\pi^{\alpha}\bar{\pi}^{\dot{\alpha}}$. This is the way the unfolded system imposes e.o.m. on primary fields: via (59), it maps $4d$ space-time fields onto an effectively $3d$ hypersurface (in the sense that $y^{\alpha}\bar{y}^{\dot{\alpha}}$ is a light-like vector). Note, however, that the role of auxiliary spinors is much more important and sophisticated. As follows directly from (11), they in fact equalize and mix translational and spin degrees of freedom. So they should not be thought of simply as coordinates on some null hypersurface in $4d$ Minkowski space.
To get a better idea of the unfolding map (59), it is instructive to consider the abelian case, where the unfolding exponent in (11) boils down to a space-time translation so that (59) can be constructed explicitly. Then a plane-wave solution to (54), (55) in Cartesian coordinates is
$A_{\alpha\dot{\alpha}}(x)=\pi_{\alpha}\bar{\mu}_{\dot{\alpha}}e^{i\pi_{\beta}\bar{\pi}_{\dot{\beta}}x^{\beta\dot{\beta}}}+c.c.,\hskip 10.00002ptF(Y|x)=i\bar{\pi}_{\dot{\alpha}}\bar{\mu}^{\dot{\alpha}}(\pi_{\alpha}y^{\alpha})^{2}e^{i\pi_{\beta}\bar{\pi}_{\dot{\beta}}(x^{\beta\dot{\beta}}+y^{\beta}\bar{y}^{\dot{\beta}})}$ (60)
with $\bar{\mu}_{\dot{\alpha}}$ being an arbitrary reference spinor, defined up to a gauge transformation $\bar{\mu}_{\dot{\alpha}}\rightarrow\bar{\mu}_{\dot{\alpha}}+const\cdot\bar{\pi}_{\dot{\alpha}}$. Putting $x=0$, one has
$\mathcal{F}(Y)=i\bar{\pi}_{\dot{\alpha}}\bar{\mu}^{\dot{\alpha}}(\pi_{\alpha}y^{\alpha})^{2}e^{i\pi_{\beta}y^{\beta}\bar{\pi}_{\dot{\beta}}\bar{y}^{\dot{\beta}}},$ (61)
which represents a plane-wave Maxwell tensor formulated purely in $Y$-terms.
In fact, one can start directly from (61) and then make use of (55) in order to fully recover $x$-dependence. It means that although we have derived the unfolded equation (55) starting from postulating (11), for (55) the expression (11) per se is nothing more than just one of many possible forms of solution. In particular, instead of (59) one can think of (55) as defining a $Y$-to-$x$ map (a similar interpretation has been proposed in [25])
$\mathcal{F}(Y)\rightarrow F(Y|x)\rightarrow F_{\alpha\alpha}(x)|_{on-shell}.$ (62)
In this picture, $\mathcal{F}(Y)$ is completely unconstrained aside from the helicity condition
[…]
This distinguishes the unfolded formulation constructed here from that of [8], where the fiber structure is tantamount to the base one. In the abelian case, the map (62) is obviously realized as
$F(Y|x)=\exp(\frac{1}{N+1}x^{\beta\dot{\beta}}\partial_{\beta}\bar{\partial}_{\dot{\beta}})\mathcal{F}_{\alpha\alpha}(y\bar{y})y^{\alpha}y^{\alpha}.$ (64)
It generates a solution to (55), and hence implicitly to Maxwell equations, for arbitrary $\mathcal{F}_{\alpha\alpha}(y\bar{y})$. This signals that relativistic dynamics in fact can be realized in terms of $Y$ without any reference to a space-time. Corresponding action principle for an arbitrary-mass integer-spin field was constructed in [26] by means of supplementing the set of $Y$ with a Lorentz-invariant proper-time coordinate serving as an evolution parameter.
The unfolding map (62) hints that the unfolded dynamics approach may provide a new way to study the problem of integrability. In particular, the question is whether it is possible to extend (64) to the non-abelian case. If so, an appropriate unfolding map will generate solutions to Yang–Mills equations in Minkowski space.
Another interesting problem is to relate the unfolded dynamics approach to twistor theory [27, 28]. Potentially, twistors may arise from treating an unfolded system as defining an unfolding map to $x$-space from some complex plane in $(Y|x)$-space, different from $x=0$ of (62) and associated with the incidence relation. Then this complex plane should be identified with the twistor space, and the corresponding unfolding map with the Penrose transform. A related discussion of these questions can be found in [29].
Conclusion
In this note, we constructed an unfolded formulation of $4d$ pure Yang–Mills theory making use of the unfolding method proposed in [18] and improved here by preliminary derivation of general relations for unfolded functions.
A natural generalization of our result would be the inclusion of charged matter. It is straightforward and should not cause any problems, as shown by the example of scalar electrodynamics [18].
Another possible direction of further research is to include supersymmetry, as well as to manifest conformal symmetry. To this end one needs to introduce the corresponding gauge 1-forms of (super)conformal gravity in addition to the Poincaré connection and to deform appropriately the unfolded equations. In its turn, this requires further non-trivial modifications of the unfolding method of [18]. In particular, it would be interesting to unfold $\mathcal{N}=4$ super-Yang–Mills theory in order to apply unfolding tools to the problems of $AdS/CFT$ and amplitudes.
Quantization of the unfolded Yang–Mills system can be performed along the lines of [12] but with necessary modifications in order to include ghosts. The first step is to build an off-shell extension of the unfolded system constructed here, which is equivalent to coupling it to external currents [30, 31].
Finally, there are two vague but potentially promising topics: the study of integrability and the derivation of twistors. Both of them seem to be related to the investigation of various unfolding maps defined by the unfolded Yang–Mills equations.
Acknowledgments
The author is grateful to A.V. Korybut, D.S. Ponomarev and M.A. Vasiliev for valuable comments and remarks.