### v3

# FIAN/TD/12-2024

Abstract
In this note, we present a novel formulation of $4d$ pure Yang-Mills theory within the unfolded framework of Vasiliev higher-spin gravity. This formulation is first-order and exhibits manifest diffeomorphism and gauge invariance. Our approach builds upon a recently proposed unfolding method, previously applied to scalar electrodynamics. Additionally, we discuss the features of various unfolding maps defined by the unfolded equations.
Introduction
Symmetry considerations serve as one of the most fundamental guiding principles in constructing theories of fundamental interactions. Higher-spin gravity theories provide prime examples, describing systems of interacting massless fields of all spins. To a large extent, these theories are determined by an infinite-dimensional higher-spin gauge symmetry [1]. For a partial overview of the relevant literature, see [2].
Maintaining control over higher-spin symmetry is of fundamental importance. The first example of higher-spin gravity was formulated through the Vasiliev equations [3, 4], a set of classical equations of motion written in the so-called unfolded form. Unfolded equations are first-order differential equations for unfolded fields – which are exterior forms – that possess manifest gauge symmetry. Thus, the unfolded dynamics approach [5, 3, 4, 6, 7, 8] provides a classical first-order formalism that ensures both manifest diffeomorphism and gauge invariance. It also enables effective control over the theory’s degrees of freedom: unfolded fields (which form an infinite spectrum in dynamical field theories) parameterize all d.o.f. of the system. In particular, these properties make the unfolded dynamics approach a promising tool for investigating $AdS/CFT$ correspondence and other dualities [9, 10, 11]. A quantization scheme for classical unfolded theories was put forward in [12] (see also [13, 14, 15, 16] for discussions of quantization of non-Lagrangian field theories).
Beyond higher-spin gravity models [3, 4, 17, 18, 19, 20], relatively few examples of unfolded formulations exist for nonlinear theories [12, 21] (see also [22] for a discussion of nonlinear unfolding in conformal geometry), thereby constraining comprehensive exploration of the approach’s full potential. The primary obstacle was the lack of a practical unfolding algorithm – especially critical for nonlinear theories. Therefore, linear models were mainly studied [23, 24, 25, 26, 27]. In [21], a novel unfolding method was put forward, that enabled the construction of an unfolded formulation of $4d$ scalar electrodynamics. In essence, it involves two steps: first postulating the form of an unfolded master-field, then deriving the corresponding unfolded equations as identities for this master-field.
In this note, we adapt and refine the method of [21] to construct an unfolded formulation of $4d$ on-shell pure Yang–Mills theory.
The note is organized as follows. First, we review basic concepts of the unfolded dynamics approach. Next, we construct the unfolded Yang–Mills equations by fixing the form of the unfolded master-fields and using certain operator relations that we derive. We then discuss the properties of the resulting unfolded system and comment on the unfolding maps it defines. In conclusion, we outline promising directions for future research.
Unfolded dynamics approach
Within the unfolded dynamics approach [5, 3, 4, 6, 7, 8], field theories are formulated via first-order differential equations
$\mathrm{d}W^{A}(x)+G^{A}(W)=0$ (1)
on unfolded fields $W^{A}(x)$, which are exterior forms on a space-time manifold $M^{d}$. Here, $A$ collectively denotes all indices of an unfolded field, $\mathrm{d}$ is the exterior derivative on $M^{d}$, and $G^{A}(W)$ is constructed from exterior products of unfolded fields (the wedge symbol is omitted throughout the paper). For each $W^{A}$ there is one and only one unfolded equation (1).
The system (1) must obey the consistency condition
$G^{B}\dfrac{\delta G^{A}}{\delta W^{B}}\equiv 0,$ (2)
which follows from $\mathrm{d}^{2}\equiv 0$ and constrains possible forms of $G^{A}$. Considering this, the unfolded equations (1) possess manifest gauge symmetry [^1] Rigorously, this holds provided the consistency is dimension-independent – a property satisfied by all known examples [8].
$\delta W^{A}=\mathrm{d}\varepsilon^{A}(x)-\varepsilon^{B}\dfrac{\delta G^{A}}{\delta W^{B}},$ (3)
with each $(n>0)$-form field $W^{A}$ giving rise to a gauge symmetry with (n1)(n-1)-form parameter $\varepsilon^{A}(x)$, while 0-form fields transform only under gauge symmetries of 1-form fields through the second term in (3).
In dynamical field theories, the spectrum of unfolded fields is infinite, as they parameterize all physical d.o.f. Typically, the space of unfolded fields admits a grading bounded from below, so that equations (1) express (perhaps, in a very complicated nonlinear way) higher-grade fields through derivatives of lower-grade ones. In addition, (1) may implicitly impose dynamical constraints on the lowest-grade fields, rendering the system on-shell. Equations (1) define what is known as a free differential algebra [28] (see also [29, 16] for discussions of the relation between the unfolded framework and various mathematical structures).
Ultimately, an unfolded system (1) describes a theory of lowest-grade fields (also referred to as primaries), possibly subject to some differential constraints (e.o.m. or whatever), while higher-grade fields constitute (infinite) towers of their covariant differential descendants. Crucially, the unfolded formulation maintains manifest diffeomorphism and gauge invariance and provides complete control over d.o.f. of the theory. These features establish unfolded dynamics approach as a powerful framework for investigating higher-spin gravity. However, these advantages extend beyond higher spins, offering potential applications to conventional field theories. In this note, our goal is to reformulate $4d$ pure Yang–Mills theory in the unfolded form (1), applying the unfolding method introduced in [21].
Yang–Mills fields and auxiliary spinors
In $sl(2,\mathbb{C})$-spinor notation, the $4d$ Yang–Mills equations together with Bianchi identities are
$\mathrm{D}_{\beta\dot{\alpha}}F^{\beta}{}_{\alpha}=0,\hskip 10.00002pt\mathrm{D}_{\alpha\dot{\beta}}\bar{F}^{\dot{\beta}}{}_{\dot{\alpha}}=0,$ (4)
where the (anti-)self-dual components of the field strength tensor are
$F_{\alpha\alpha}:=\frac{\partial}{\partial x^{\alpha\dot{\beta}}}A_{\alpha}{}^{\dot{\beta}}-i[A_{\alpha\dot{\beta}},A_{\alpha}{}^{\dot{\beta}}],\hskip 10.00002pt\bar{F}_{\dot{\alpha}\dot{\alpha}}:=\frac{\partial}{\partial x^{\beta\dot{\alpha}}}A^{\beta}{}_{\dot{\alpha}}-i[A_{\beta\dot{\alpha}},A^{\beta}{}_{\dot{\alpha}}$], (5)
and the covariant derivative is
$\mathrm{D}_{\alpha\dot{\alpha}}:=\frac{\partial}{\partial x^{\alpha\dot{\alpha}}}-i[A_{\alpha\dot{\alpha}},\bullet],$ (6)
satisfying
$\mathrm{D}_{\alpha\dot{\alpha}},\mathrm{D}_{\beta\dot{\beta}}]=-i\epsilon_{\alpha\beta}\bar{F}_{\dot{\alpha}\dot{\beta}}-i\epsilon_{\dot{\alpha}\dot{\beta}}F_{\alpha\beta}.$ (7)
The antisymmetric spinor metric
$\epsilon_{\alpha\beta}=\epsilon_{\dot{\alpha}\dot{\beta}}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),\quad\epsilon^{\alpha\beta}=\epsilon^{\dot{\alpha}\dot{\beta}}=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right)$ (8)
raises/lowers spinor indices via
$v_{\alpha}=\epsilon_{\beta\alpha}v^{\beta},\quad v^{\alpha}=\epsilon^{\alpha\beta}v_{\beta},\quad\bar{v}_{\dot{\alpha}}=\epsilon_{\dot{\beta}\dot{\alpha}}\bar{v}^{\dot{\beta}},\quad\bar{v}^{\dot{\alpha}}=\epsilon^{\dot{\alpha}\dot{\beta}}\bar{v}_{\dot{\beta}}.$ (9)
For multispinors, indices denoted with the same letter are either contracted or symmetrized, depending on their relative positions, (similarly for dotted indices)
$T_{\alpha\alpha}:=T_{(\alpha_{1}\alpha_{2})},\quad T_{\alpha}{}^{\alpha}:=\epsilon^{\alpha\beta}T_{\alpha\beta}.$ (10)
In order to unfold Yang–Mills theory, one has to introduce, on top of primaries $F_{\alpha\alpha}$ and $\bar{F}_{\dot{\alpha}\dot{\alpha}}$, the infinite towers of all their differential on-shell descendants. This can be conveniently performed by means of auxiliary commuting spinors [^2] In $4d$ Vasiliev equations [3, 4], these spinors become generators of an associative higher-spin gauge algebra through a specific non-commutative star product defined on them. $Y=(y^{\alpha},\bar{y}^{\dot{\alpha}})$, then the whole towers get packed into unfolded Yang–Mills master-fields, which we postulate to be of the form
$F(Y|x)=e^{\mathrm{D}y\bar{y}}F_{\alpha\alpha}(x)y^{\alpha}y^{\alpha}e^{-\mathrm{D}y\bar{y}},\quad\bar{F}(Y|x)=e^{\mathrm{D}y\bar{y}}\bar{F}_{\dot{\alpha}\dot{\alpha}}(x)\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}e^{-\mathrm{D}y\bar{y}},$ (11)
where right-acting derivatives are contracted with spinors as
$\mathrm{D}y\bar{y}:=y^{\alpha}\bar{y}^{\dot{\alpha}}\mathrm{D}_{\alpha\dot{\alpha}}.$ (12)
Unfolded master-fields (11) contain primary Yang-Mills tensors as (anti-)holomorphic in $Y$ components
$F_{\alpha\alpha}(x)y^{\alpha}y^{\alpha}=F(Y|x)|_{\bar{y}=0},\quad\bar{F}_{\dot{\alpha}\dot{\alpha}}(x)\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}=\bar{F}(Y|x)|_{y=0},$ (13)
together with an infinite sequence of their fully symmetrized traceless covariant derivatives of all orders. This constitutes the set of all independent covariant descendants of the primary Yang–Mills tensor, since antisymmetrizations and contractions are determined by (4) and (7). By construction, unfolded master-fields (11) inherit the adjoint representation of the gauge algebra of primary tensors.
The aforementioned grading on the space of unfolded fields can be introduced in terms of the spinor Euler operators
$N:=y^{\alpha}\partial_{\alpha},\quad\bar{N}:=\bar{y}^{\dot{\alpha}}\bar{\partial}_{\dot{\alpha}},$ (14)
where $\partial_{\alpha}$ and $\bar{\partial}_{\dot{\alpha}}$ are $y^{\alpha}$- and $\bar{y}^{\dot{\alpha}}$-derivatives. From their definitions (11), master-fields obey
$(N-\bar{N})F=2F,\quad(N-\bar{N})\bar{F}=-2\bar{F},$ (15)
which corresponds to helicities $\pm 1$ (in higher-spin gravity, strength tensors with $|N-\bar{N}|F=2sF$ describe spin-$s$ massless fields) and, further,
$\mathrm{D}y\bar{y}F=\bar{N}F,\quad\mathrm{D}y\bar{y}\bar{F}=N\bar{F}.$ (16)
Then a grading operator on the space of unfolded fields can be defined as
$\mathcal{G}:=\frac{1}{2}(N+\bar{N}-2).$ (17)
It decomposes the space of master-fields $F$ into a direct sum of its eigenspaces (the same applies to $\bar{F}$)
$F(Y|x)=\sum_{n=0}^{\infty}F^{(n)}(Y|x),\quad\mathcal{G}F^{(n)}=nF^{(n)},$ (18)
so that the primary has 0-grade, while $n$-grade unfolded fields represent its covariant derivatives of $n$-th order, as follows from (11).
To formulate the unfolded equations (1), one needs to express the derivatives of the unfolded fields in algebraic terms. In our particular case, the task is to express $\mathrm{D}_{\alpha\dot{\beta}}F$ for $F$ (11) in terms of $Y$ and $\partial/\partial Y$ acting on $F$ and $\bar{F}$ (and the same for $\mathrm{D}_{\alpha\dot{\beta}}\bar{F}$). Before we start processing (11), let us work out some general operator relations that simplify the analysis implied by the method of [21]. Combining a commutator formula
$\hat{\mathrm{A}},e^{\hat{\mathrm{D}}}]=\int_{0}^{1}dte^{t\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-t\hat{\mathrm{D}}}e^{\hat{\mathrm{D}}}$ (19)
with an Euler-operator representation of a homotopy integral
$\int_{0}^{1}dtt^{k}F(tz)=\frac{1}{z\frac{\partial}{\partial z}+1+k}F(z),$ (20)
one obtains
[…]
where the inverse Euler operator is understood (when $F(0)=0$) as
[…]
Applying this formula to an operator of the form
$\hat{\mathrm{B}}=e^{\hat{\mathrm{D}}}\hat{\mathrm{C}}e^{-\hat{\mathrm{D}}}$ (23)
yields
$\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}}e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{D}}]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}$. (24)
Next, considering the case $\hat{\mathrm{C}}=\hat{\mathrm{D}}$, from (24) one gets
[…]
so that (24) can be equivalently rewritten as
$\hat{\mathrm{A}},\hat{\mathrm{B}}]=[(\frac{1}{N_{\hat{\mathrm{D}}}(N_{\hat{\mathrm{D}}}-1)}e^{\hat{\mathrm{D}}}[\hat{\mathrm{D}},[\hat{\mathrm{A}},\hat{\mathrm{D}}]]e^{-\hat{\mathrm{D}}}),\hat{\mathrm{B}}]+[[\hat{\mathrm{A}},\hat{\mathrm{D}}],\hat{\mathrm{B}}]+e^{\hat{\mathrm{D}}}[\hat{\mathrm{A}},\hat{\mathrm{C}}]e^{-\hat{\mathrm{D}}}$. (26)
For convenience, we denote an "unfolding map" for an arbitrary field $\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)$ taking values in the adjoint representation as
$\ll\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)\gg:=e^{\mathrm{D}y\bar{y}}\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)e^{-\mathrm{D}y\bar{y}},$ (27)
so that, in particular,
$F(Y|x)=\ll F_{\alpha\alpha}(x)y^{\alpha}y^{\alpha}\gg,\quad\bar{F}(Y|x)=\ll\bar{F}_{\dot{\alpha}\dot{\alpha}}(x)\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg.$ (28)
Then from (26) one has (here $C\equiv\mathrm{C}_{\alpha(n),\dot{\beta}(m)}(Y|x)$ )
$\ll\partial_{\mu}\mathrm{C}\gg=(\partial_{\mu}-\mathrm{D}_{\mu\dot{\mu}}\bar{y}^{\dot{\mu}})\ll\mathrm{C}\gg+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\ll\mathrm{C}\gg],$ (29)
$\ll\bar{\partial}_{\dot{\mu}}\mathrm{C}\gg=(\bar{\partial}_{\dot{\mu}}-\mathrm{D}_{\mu\dot{\mu}}y^{\mu})\ll\mathrm{C}\gg+i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+1)(\bar{N}+2)}F,\ll\mathrm{C}\gg].$ (30)
From here on, all derivatives inside the angle brackets always act only on an expression within brackets and never differentiate unfolding exponents of (27). The square brackets stand for the commutator in the gauge Lie algebra and the Euler operators of (26) are expressed in terms of the spinor Euler operators (14). Master-fields $F$ and $\bar{F}$ arise in (29)-(30) through (7) without assuming Yang–Mills equations (4).
Now, let us get down to solving the problem. Direct application of (24) to $\mathrm{D}_{\mu\dot{\mu}}F(Y|x)$ gives
$\mathrm{D}_{\mu\dot{\mu}}F=[\frac{1}{N}\ll[\mathrm{D}_{\mu\dot{\mu}},\mathrm{D}y\bar{y}]\gg,F]+\ll\mathrm{D}_{\mu\dot{\mu}}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg.$ (31)
The task is to eliminate all covariant derivatives and angle brackets on the r.h.s. by re-expressing them in terms of unfolded master-fields $F$ and $\bar{F}$ acted on by $Y$’s and $\frac{\partial}{\partial Y}$’s. To accomplish this, we have the following tools at our disposal: Yang–Mills equations (4) together with (7), the relations (29) and (30), the Schouten identities for spinors and the Jacobi identity of the gauge Lie algebra.
Applying (7) to the first term on the r.h.s. of (31) and (4) (plus Schouten identities) to the second one, one has
$\mathrm{D}_{\mu\dot{\mu}}F=\frac{i}{2}\bar{y}_{\dot{\mu}}[\frac{1}{N}\ll\partial_{\mu}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg,F]+\frac{i}{2}y_{\mu}[\frac{1}{\bar{N}}\ll\bar{\partial}_{\dot{\mu}}\bar{F}_{\dot{\alpha}\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg,F]+\frac{1}{3}\ll\partial_{\mu}\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg.$ (32)
First, we process the first term on the r.h.s. (the second term will be resolved by conjugation). Applying (29) to the problematic factor yields
$\ll\partial_{\mu}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\partial_{\mu}F-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F].$ (33)
Therefore, one needs to process $\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F$. Contracting (32) with $\bar{y}^{\dot{\mu}}$ gives
$\mathrm{D}_{\mu\dot{\mu}}\bar{y}^{\dot{\mu}}F=iy_{\mu}[\frac{1}{\bar{N}-1}\bar{F},F]+\frac{1}{3}\ll\partial_{\mu}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg.$ (34)
On the other hand, from (16) and (29) one finds
$\ll\partial_{\mu}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\partial_{\mu}\bar{N}F-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{N}F+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\bar{N}F].$ (35)
Combining (34) and (35), one obtains after rearranging the Euler-operator ratios
$\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}}F=\frac{\bar{N}}{N+1}\partial_{\mu}F+iy_{\mu}\frac{2}{N+2}[\frac{1}{N+2}\bar{F},F]+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F].$ (36)
Thus one finds for (33)
$\ll\partial_{\mu}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\frac{2}{N+1}(\partial_{\mu}F-iy_{\mu}[\frac{1}{N+2}\bar{F},F]).$ (37)
Conjugation gives
$\ll\bar{\partial}_{\dot{\mu}}\bar{F}_{\dot{\alpha}\dot{\alpha}}\bar{y}^{\dot{\alpha}}\bar{y}^{\dot{\alpha}}\gg=\frac{2}{\bar{N}+1}(\bar{\partial}_{\dot{\mu}}\bar{F}-i\bar{y}_{\dot{\mu}}[\frac{1}{\bar{N}+1}F,\bar{F}]).$ (38)
These bring first two terms on the r.h.s. of (32) to the admissible form.
Now we process the last term in (32). By virtue of (29) one has
$\ll\partial_{\mu}\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=(\partial_{\mu}-\mathrm{D}_{\mu\dot{\alpha}}\bar{y}^{\dot{\alpha}})\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg+$
$+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg].$ (39)
Since obviously
$\ll\bar{\partial}_{\dot{\mu}}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=0,$ (40)
one has from (30)
$\mathrm{D}_{\alpha\dot{\mu}}y^{\alpha}F=\bar{\partial}_{\dot{\mu}}F+i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+1)(\bar{N}+2)}F,F].$ (41)
Using this result and contracting (32) with $y^{\mu}$ yields
$\ll\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\bar{\partial}_{\dot{\mu}}F-i\bar{y}_{\dot{\mu}}[\frac{1}{(\bar{N}+2)}F,F],$ (42)
which together with (36) turns (39) to
$\ll\partial_{\mu}\bar{\partial}_{\dot{\mu}}\mathrm{D}y\bar{y}F_{\alpha\alpha}y^{\alpha}y^{\alpha}\gg=\partial_{\mu}\bar{\partial}_{\dot{\mu}}F-i\bar{y}_{\dot{\mu}}\partial_{\mu}[\frac{1}{\bar{N}+2}F,F]+i[\frac{1}{(N+1)(N+2)}\bar{F},y_{\mu}\bar{\partial}_{\dot{\mu}}F]+$
$+y_{\mu}\bar{y}_{\dot{\mu}}[\frac{1}{(N+1)(N+2)}\bar{F},[\frac{1}{\bar{N}+2}F,F]]+\mathrm{D}_{\mu\dot{\mu}}F-\frac{\bar{N}+1}{N+1}\partial_{\mu}\bar{\partial}_{\dot{\mu}}F$-
$-2iy_{\mu}\bar{\partial}_{\dot{\mu}}\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F]-iy_{\mu}\bar{\partial}_{\dot{\mu}}[\frac{1}{(N+1)(N+2)}\bar{F},F]+$
$+i\bar{y}_{\dot{\mu}}[\frac{1}{\bar{N}+1}(\frac{\bar{N}}{N+1}\partial_{\mu}F+iy_{\mu}\frac{2}{N+2}[\frac{1}{N+2}\bar{F},F]+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F]),F]+$
$+i\bar{y}_{\dot{\mu}}[\frac{1}{N}F,(\frac{\bar{N}}{N+1}\partial_{\mu}F+iy_{\mu}\frac{2}{N+2}[\frac{1}{N+2}\bar{F},F]+iy_{\mu}[\frac{1}{(N+1)(N+2)}\bar{F},F])].$ (43)
Substituting (37), (38) and (43) into (32) and combining like terms using the gauge algebra Jacobi identity, one finally expresses the covariant derivative of $F$ in $Y$-terms
$\mathrm{D}_{\mu\dot{\mu}}F=\frac{1}{N+1}\partial_{\mu}\bar{\partial}_{\dot{\mu}}F+iN[\frac{1}{N(N+1)}\bar{y}_{\dot{\mu}}\partial_{\mu}F,\frac{1}{N}F]-iy_{\mu}\bar{\partial}_{\dot{\mu}}\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F]+$
$+[\frac{i}{(N+1)(N+2)}y_{\mu}\bar{\partial}_{\dot{\mu}}\bar{F},F]+\frac{1}{2}y_{\mu}\bar{y}_{\dot{\mu}}[\frac{N+3}{(N+1)(N+2)}[\frac{1}{N+2}\bar{F},F],F]+$
$+\frac{3}{2}y_{\mu}\bar{y}_{\dot{\mu}}[\frac{1}{(N+1)(N+2)}[\frac{1}{N}F,\bar{F}],F]+y_{\mu}\bar{y}_{\dot{\mu}}[\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F],\frac{1}{N}F].$ (44)
Poincaré symmetry and diffeomorphism-invariance
The manifestly Poincaré-invariant relation (44) is written in Cartesian coordinates, together with (6) and (11), since they involve $\frac{\partial}{\partial x^{\mu\dot{\mu}}}$ inside of $\mathrm{D}_{\mu\dot{\mu}}$. The unfolded dynamics approach requires manifest coordinate-independence, which is ensured by formulating equations in terms of exterior forms. To attain this, we switch to the fiber space picture: we claim that $F(Y|x)$ and $\bar{F}(Y|x)$ are now 0-forms on the Minkowski base manifold with some local coordinates $x^{\underline{n}}$, while spinor variables $\{y^{\alpha},\bar{y}^{\dot{\alpha}}\}$ are coordinates in the fiber. This requires appropriate generalization of $\mathrm{D}_{\mu\dot{\mu}}$ in (44).
According to the ideology of unfolding, global Poincaré symmetry of a relativistic theory should arise in terms of the unique general formula (3). This is achieved by introducing a non-dynamical 1-form $\Omega(x)$, which takes values in Lie algebra $iso(1,3)$
$\Omega=e^{\alpha\dot{\beta}}P_{\alpha\dot{\beta}}+\omega^{\alpha\alpha}M_{\alpha\alpha}+\bar{\omega}^{\dot{\alpha}\dot{\alpha}}\bar{M}_{\dot{\alpha}\dot{\alpha}},$ (45)
with $P_{\alpha\dot{\alpha}}$, $M_{\alpha\alpha}$ and $\bar{M}_{\dot{\alpha}\dot{\alpha}}$ being generators of translations and rotations of $\mathbb{R}^{1,3}$, and $e^{\alpha\dot{\beta}}$ and $\omega^{\alpha\alpha}$ ($\bar{\omega}^{\dot{\alpha}\dot{\alpha}}$) being 1-forms of a vierbein and a Lorentz connection.
$\Omega$ is subjected to the flatness condition (square brackets stand for the $iso(1,3)$-commutator)
$\mathrm{d}\Omega+\frac{1}{2}[\Omega,\Omega]=0,$ (46)
so that the corresponding gauge symmetry (3) is
$\delta\Omega=\mathrm{d}\varepsilon(x)+[\Omega,\varepsilon]$ (47)
and describes an infinite-dimensional freedom in switching between all possible local coordinates on $\mathbb{R}^{1,3}$. This boils down to 10-dimensional global Poincaré symmetry after fixing some particular solution $\Omega_{0}$ and restricting to those residual $\varepsilon(x)$ which leave it invariant
$\mathrm{d}\varepsilon_{0}+[\Omega_{0},\varepsilon_{0}]=0.$ (48)
The simplest non-degenerate global solution to (46) is provided by Cartesian coordinates
$e_{\underline{m}}{}^{\alpha\dot{\beta}}=(\bar{\sigma}_{\underline{m}})^{\dot{\beta}\alpha},\quad\omega_{\underline{m}}{}^{\alpha\alpha}=0,\quad\bar{\omega}_{\underline{m}}{}^{\dot{\alpha}\dot{\alpha}}=0,$ (49)
with global symmetries (48) parameterized by $x$-independent $\xi^{\alpha\dot{\beta}}$, $\xi^{\alpha\alpha}$ and $\bar{\xi}^{\dot{\alpha}\dot{\alpha}}$
$\varepsilon_{0}^{\alpha\dot{\beta}}=\xi^{\alpha\dot{\beta}}+\xi^{\alpha}\text{}_{\gamma}(\bar{\sigma}_{\underline{m}})^{\dot{\beta}\gamma}x^{\underline{m}}+\bar{\xi}^{\dot{\beta}}\text{}_{\dot{\gamma}}(\bar{\sigma}_{\underline{m}})^{\dot{\gamma}\alpha}x^{\underline{m}},\quad\varepsilon_{0}^{\alpha\alpha}=\xi^{\alpha\alpha},\quad\bar{\varepsilon}_{0}^{\dot{\alpha}\dot{\alpha}}=\bar{\xi}^{\dot{\alpha}\dot{\alpha}}.$ (50)
Analogously, in order to realize the Yang–Mills gauge symmetry via (3), one introduces a 1-form $A(x)$, with $A_{\alpha\dot{\alpha}}(x)$ being its expansion in the vierbein
$A(x)=e^{\alpha\dot{\alpha}}A_{\alpha\dot{\alpha}}.$ (51)
Now an appropriate coordinate-independent generalization of $\mathrm{D}_{\mu\dot{\mu}}$ is a 1-form operator $\mathrm{D}$, supplemented by Lorentz-connection terms rotating fiber coordinates $Y$,
$\mathrm{D}:=\mathrm{d}+\omega^{\alpha\alpha}y_{\alpha}\partial_{\alpha}+\bar{\omega}^{\dot{\alpha}\dot{\alpha}}\bar{y}_{\dot{\alpha}}\bar{\partial}_{\dot{\alpha}}-i[A,\bullet].$ (52)
In Cartesian coordinates (49), this indeed boils down to
$\mathrm{D}=\mathrm{d}x^{\mu\dot{\mu}}\mathrm{D}_{\mu\dot{\mu}}.$ (53)
Unfolded Yang–Mills equations and unfolding maps
Now we are ready to write down an unfolded system for Yang–Mills theory, which is the main result of the note. Contracting (5) and (44) with the vierbeins yields
$\mathrm{d}A+[A,A]=\frac{1}{4}e^{\alpha}{}_{\dot{\beta}}e^{\alpha\dot{\beta}}\partial_{\alpha}\partial_{\alpha}F|_{\bar{y}=0}+\frac{1}{4}e_{\beta}{}^{\dot{\alpha}}e^{\beta\dot{\alpha}}\bar{\partial}_{\dot{\alpha}}\bar{\partial}_{\dot{\alpha}}\bar{F}|_{y=0},$ (54)
$\mathrm{D}F=\frac{1}{N+1}e\partial\bar{\partial}F+iN[\frac{1}{N(N+1)}e\partial\bar{y}F,\frac{1}{N}F]-iey\bar{\partial}\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F]+$
$+[\frac{i}{(N+1)(N+2)}ey\bar{\partial}\bar{F},F]+\frac{1}{2}ey\bar{y}[\frac{N+3}{(N+1)(N+2)}[\frac{1}{N+2}\bar{F},F],F]+$
$+\frac{3}{2}ey\bar{y}[\frac{1}{(N+1)(N+2)}[\frac{1}{N}F,\bar{F}],F]+ey\bar{y}[\frac{1}{N+2}[\frac{1}{N+2}\bar{F},F],\frac{1}{N}F],$ (55)
plus a conjugate equation for $\bar{F}$ resulting from exchanging barred and unbarred objects in (55). Here
$e\partial\bar{\partial}:=e^{\alpha\dot{\beta}}\partial_{\alpha}\bar{\partial}_{\dot{\beta}},\quad e\partial\bar{y}:=e^{\alpha\dot{\beta}}\partial_{\alpha}\bar{y}_{\dot{\beta}},\quad ey\bar{\partial}:=e^{\alpha\dot{\beta}}y_{\alpha}\bar{\partial}_{\dot{\beta}}\quad ey\bar{y}:=e^{\alpha\dot{\beta}}y_{\alpha}\bar{y}_{\dot{\beta}}.$ (56)
A full spectrum of unfolded fields consists of a 1-form $\Omega$ describing Minkowski background, a 1-form of the gauge potential $A$ and the 0-forms of master-fields $F(Y|x)$ and $\bar{F}(Y|x)$ encoding the Yang–Mills tensor together with an infinite tower of its covariant derivatives. The corresponding unfolded equations are (46), (54) and (55) plus a conjugate for $\bar{F}$. The formulation is manifestly diffeomorphism-invariant. 1-forms $\Omega$ and $A$ give rise to two manifest symmetries in accordance with (3): the global (after fixing $\Omega$) Poincaré one and the local Yang–Mills one. The Yang–Mills symmetry is realized as
$\delta A(x)=\mathrm{D}\varepsilon(x),\quad\delta F(Y|x)=i[\varepsilon(x),F(Y|x)],\quad\delta\bar{F}(Y|x)=i[\varepsilon(x),\bar{F}(Y|x)].$ (57)
By construction, the unfolded system is consistent, provided unfolded 0-form fields obey the helicity constraint (15). At the same time, a direct check of (2) seems hardly executable due to the complexity of the equations. Note that the system is at most cubic in master-field 0-forms, which is surprising on its own. One of the possible forms of solution to the system is (11) with primary fields subjected to (4). Let us quickly derive this directly from (55).
Master-fields $F$ and $\bar{F}$ are assumed to be analytical in $Y$ and obey (15). For the sake of brevity, here and below we consider only an anti-selfdual component $F$, but everything applies to $\bar{F}$ as well. Acting on (55) with $y^{\beta}\bar{y}^{\dot{\beta}}\frac{\delta}{\delta e^{\beta\dot{\beta}}}$ produces (16), whose solution, accounting for (15), is (11). On the other hand, acting on (55) with $\frac{1}{2}\partial_{\mu}\partial^{\beta}\frac{\delta}{\delta e^{\beta\dot{\mu}}}$ and putting $\bar{y}^{\dot{\alpha}}=0$ yields, accounting for (15) again,
$\mathrm{D}_{\beta\dot{\mu}}F^{\beta}{}_{\mu}=0.$ (58)
Thus, the unfolded system presented above indeed provides a consistent manifestly diffeomorphism- and gauge-invariant first-order formulation of $4d$ Yang-Mills theory. Unfolded master-fields $F(Y|x)$ and $\bar{F}(Y|x)$ encode all on-shell d.o.f. of the Yang-Mills tensor as expansions in auxiliary spinors $Y$.
The system allows for two obvious reductions. The first one is the anti-selfdual case $\bar{F}(Y|x)=0$ (or the selfdual $F(Y|X)=0$), then only first two terms on the r.h.s. of (55) survive. In different (tensor) terms, this was presented in [30]. The second one is the abelian case with all commutators vanishing, then only the first term on the r.h.s. of (55) remains.
One can think of the equation (55) as defining an unfolding map from $x$-space to $Y$-space
$F_{\alpha\alpha}(x)|_{on-shell}\rightarrow F(Y|x)\rightarrow\mathcal{F}(Y):=F(Y|x=0).$ (59)
Then (11) explicitly realizes the first arrow in (59). The field $\mathcal{F}(Y)$ carries precisely the same information as on-shell $F_{\alpha\alpha}$ does. In a sense, spinors $Y=\{y^{\alpha};\bar{y}^{\dot{\alpha}}\}$ effectively replace space-time coordinates $x^{\alpha\dot{\alpha}}$ for on-shell configurations, hence being conjugate to spinor-helicity variables that resolve light-like momenta $p^{\alpha\dot{\alpha}}=\pi^{\alpha}\bar{\pi}^{\dot{\alpha}}$. This is the way the unfolded system imposes e.o.m. on primary fields: via (59), it maps $4d$ space-time fields onto an effectively $3d$ hypersurface (in the sense that $y^{\alpha}\bar{y}^{\dot{\alpha}}$ is a light-like vector). Note however, that the role of auxiliary spinors is much more important and sophisticated. As follows directly from (11), they in fact equalize and mix translational and spin degrees of freedom. So they should not be thought of simply as coordinates on some null hypersurface in $4d$ Minkowski space.
To get a better idea of the unfolding map (59), it is instructive to consider the abelian case, where the unfolding exponent in (11) boils down to a space-time translation so that (59) can be constructed explicitly. Then a plane-wave solution to (54), (55) in Cartesian coordinates is
$A_{\alpha\dot{\alpha}}(x)=\pi_{\alpha}\bar{\mu}_{\dot{\alpha}}e^{i\pi_{\beta}\bar{\pi}_{\dot{\beta}}x^{\beta\dot{\beta}}}+c.c.,\hskip 10.00002ptF(Y|x)=i\bar{\pi}_{\dot{\alpha}}\bar{\mu}^{\dot{\alpha}}(\pi_{\alpha}y^{\alpha})^{2}e^{i\pi_{\beta}\bar{\pi}_{\dot{\beta}}(x^{\beta\dot{\beta}}+y^{\beta}\bar{y}^{\dot{\beta}})}$ (60)
with $\bar{\mu}_{\dot{\alpha}}$ being an arbitrary reference spinor, defined up to a gauge transformation $\bar{\mu}_{\dot{\alpha}}\rightarrow\bar{\mu}_{\dot{\alpha}}+const\cdot\bar{\pi}_{\dot{\alpha}}$. Putting $x=0$, one has
$\mathcal{F}(Y)=i\bar{\pi}_{\dot{\alpha}}\bar{\mu}^{\dot{\alpha}}(\pi_{\alpha}y^{\alpha})^{2}e^{i\pi_{\beta}y^{\beta}\bar{\pi}_{\dot{\beta}}\bar{y}^{\dot{\beta}}},$ (61)
which represents a plane-wave Maxwell tensor formulated purely in $Y$-terms.
In fact, one can start directly from (61) and then make use of (55) in order to fully recover $x$-dependence. This implies that although we have derived the unfolded equation (55) starting from postulating (11), for (55) the expression (11) per se is nothing more than just one of many possible forms of solution. In particular, instead of (59) one can think of (55) as defining a $Y$-to-$x$ map (a similar interpretation has been proposed in [31])
$\mathcal{F}(Y)\rightarrow F(Y|x)\rightarrow F_{\alpha\alpha}(x)|_{on-shell}.$ (62)
In this picture, $\mathcal{F}(Y)$ is completely unconstrained aside from the helicity condition
$(N-\bar{N})\mathcal{F}(Y)=2\mathcal{F}(Y)\hskip 10.00002pt\Rightarrow\hskip 10.00002pt\mathcal{F}(Y)=\mathcal{F}_{\alpha\alpha}(y\bar{y})y^{\alpha}y^{\alpha}.$ (63)
This distinguishes the unfolded formulation constructed here from that of [8], where the fiber structure is tantamount to the base one. In the abelian case, the map (62) is obviously realized as
$F(Y|x)=\exp(\frac{1}{N+1}x^{\beta\dot{\beta}}\partial_{\beta}\bar{\partial}_{\dot{\beta}})\mathcal{F}_{\alpha\alpha}(y\bar{y})y^{\alpha}y^{\alpha}.$ (64)
It generates a solution to (55), and hence implicitly to Maxwell equations, for arbitrary $\mathcal{F}_{\alpha\alpha}(y\bar{y})$. This indicates that relativistic dynamics in fact can be realized in terms of $Y$ without any reference to a space-time. The corresponding action principle for an arbitrary-mass integer-spin field was constructed in [32] by means of supplementing the set of $Y$ with a Lorentz-invariant proper-time coordinate serving as an evolution parameter.
From this perspective, Yang–Mills theory is characterized by the unconstrained field $\mathcal{F}(Y)=\mathcal{F}_{\alpha\alpha}(y\bar{y})y^{\alpha}y^{\alpha}$ and its conjugate $\bar{\mathcal{F}}$. These field form representations of both the gauge and Poincaré algebras, as defined by the unfolded system via the general formula (3). At first glance, this might suggest the theory is already solved, since no equations constrain $\mathcal{F}$ and $\bar{\mathcal{F}}$. However, the key complication preserving the theory’s non-triviality lies in the nonlinear realization of Poincaré translations in $\mathcal{F}$ and $\bar{\mathcal{F}}$, as evident from (55). Considering this, the unfolding map (62) suggests that the unfolded dynamics approach may provide a new way to study the problem of integrability of classical e.o.m. and constructing exact solutions. In particular, the question is whether it is possible to extend (64) to the non-abelian case. If so, an appropriate unfolding map will generate solutions to Yang–Mills equations in Minkowski space. Additionally, the unfolded framework provides a systematic method for identifying theory invariants: all gauge-invariant conserved charges are classified through the cohomology of a certain operator determined by the unfolded equations [8].
Another interesting problem is to relate the unfolded dynamics approach to twistor theory [33, 34]. Potentially, twistors may arise from treating an unfolded system as defining an unfolding map to $x$-space from some complex plane in $(Y|x)$-space, different from $x=0$ of (62) and associated with the incidence relation. Then this complex plane should be identified with the twistor space, and the corresponding unfolding map with the Penrose transform. A related discussion of these questions can be found in [35].
Conclusion
In this note, we constructed an unfolded formulation of $4d$ pure Yang–Mills theory making use of the unfolding method proposed in [21] and improved here by the preliminary derivation of general relations for unfolded functions.
A natural generalization of our result would be the inclusion of charged matter. It is straightforward and should not present any difficulties, as shown by the example of scalar electrodynamics [21].
Another possible direction of further research is to include supersymmetry, as well as to manifest conformal symmetry. To this end one needs to introduce the corresponding gauge 1-forms of (super)conformal gravity in addition to the Poincaré connection and to deform appropriately the unfolded equations. In its turn, this requires further non-trivial modifications of the unfolding method of [21]. In particular, it would be interesting to unfold $\mathcal{N}=4$ super-Yang–Mills theory in order to apply unfolding tools to the problems of $AdS/CFT$ and amplitudes. Another interesting problem is to unfold Einstein equations.
Quantization of the unfolded Yang–Mills system can be performed along the lines of [12] but with necessary modifications in order to include ghosts. The first step is to build an off-shell extension of the unfolded system constructed here, which is equivalent to coupling it to external currents [36, 37]. After that, this off-shell system can be elevated to an unfolded Schwinger–Dyson system for correlation functions of quantum unfolded fields.
Finally, there are two vague but potentially promising topics: the study of integrability of classical e.o.m. and the derivation of twistors. Both of them seem to be related to the investigation of various unfolding maps defined by the unfolded Yang–Mills equations.
Acknowledgments
The author is grateful to A.V. Korybut, D.S. Ponomarev and M.A. Vasiliev for valuable comments and remarks.